OpenStudy (anonymous):

Maria has three children. There is two years age difference between each child. The total ages of all three children is 36 years. Rosa is the youngest child. How old is rosa?

OpenStudy (dan815):

u have 3 ages and they are 2 apart so thats 6 years + 3 ages of same ppl 3x+6=36

OpenStudy (dan815):

or u wanna write one by one u can do that too lets say x is the age of rosa the younger x+2 = rosas sister x+2+2 = rosas brother their ages together add up to 36 X+X+2+X+2+2 = 36 3X+6=36 X=10 SO ROSA MUST BE 10

OpenStudy (anonymous):

i have a friend called rosa. she is definitely not 10.

OpenStudy (anonymous):

false accusations. :P

OpenStudy (dan815):

rosa is a *****

OpenStudy (anonymous):

HAHAHAHAHAHAHA

OpenStudy (anonymous):

medal u for that

OpenStudy (dan815):

give it to me! i need my lvl

OpenStudy (anonymous):

Thank you

OpenStudy (dan815):

daryl select best comment gimme medal :D

OpenStudy (anonymous):

greedy

OpenStudy (dan815):

YESS I got it 73:)

OpenStudy (anonymous):

give it to me.

OpenStudy (dan815):

ty

OpenStudy (anonymous):

u can always undo

OpenStudy (anonymous):

Did I do it?

OpenStudy (dan815):

noo dont u dare

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Did I give Dan a medal?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay :)thanks! Imma leave now.

OpenStudy (dan815):

cya